# Will Trump Win in 2024? Predicting the US Presidential Election via Multi-step Reasoning with Large Language Models

### v1

@@ Abstract
Can Large Language Models (LLMs) accurately predict election outcomes? While LLMs have demonstrated impressive performance in various domains, including healthcare, legal analysis, and creative tasks, their ability to forecast elections remains unknown. Election prediction poses unique challenges, such as limited voter-level data, rapidly changing political landscapes, and the need to model complex human behavior. To address these challenges, we introduce a multi-step reasoning framework designed for political analysis. Our approach is validated on real-world data from the American National Election Studies (ANES) 2016 and 2020, as well as synthetic personas generated by the leading machine learning framework, offering scalable datasets for voter behavior modeling. To capture temporal dynamics, we incorporate candidates’ policy positions and biographical details, ensuring that the model adapts to evolving political contexts. Drawing on Chain of Thought prompting, our multi-step reasoning pipeline systematically integrates demographic, ideological, and time-dependent factors, enhancing the model’s predictive power. Also, we apply our framework to predict the outcome of the 2024 U.S. presidential election in advance, demonstrating the adaptability of LLMs to unseen political data.
Will Trump Win in 2024? Predicting the US Presidential Election via Multi-step Reasoning with Large Language Models
Chenxiao Yu1, Zhaotian Weng1, Zheng Li2, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Carnegie Mellon University {cyu96374, wengzhao, yzhao010}@usc.edu, winston@arimadata.com, xiyanghu@cmu.edu Corresponding authors
Important Notice Ongoing Work: This research is ongoing as of October 21, 2024. Research Integrity: This research is conducted independently without any external funding or support. Content Warning: This paper may contain some offensive content generated by LLMs.
@@ 3.1.2 Synthetic Data for the US Population
Regarding the nationwide 2024 prediction discussed in Section 4, we will run a comprehensive simulation across all states, excluding Rhode Island (RI) and Connecticut (CT) due to missing data. This allows us to better evaluate the LLM’s ability to generate results on previously unseen datasets.
@@ 3.3.2 Evaluations on Synthetic Data for the 2020 US Population
Figure 6: LLM’s prediction for the 2024 election. The predicted probability for each state is calculated as $P=\frac{R}{R+D}$, where $R$ and $D$ represent the total votes for the Republican and Democratic candidates, respectively. States shaded in red ( ) indicate a predicted Republican win, while states shaded in blue ( ) indicate a predicted Democratic win. It is worth noting that for Nebraska (NE), we did not use the congressional district method to calculate electoral votes but instead applied the winner-takes-all method, which is consistent with other states. Additionally, due to data limitations, we excluded Connecticut (CT, 7 electoral votes) and Rhode Island (RI, 4 electoral votes), denoted in from the voting simulation, resulting in a total of 527 electoral votes (538 - 11). Based on the LLM’s predictions, Donald Trump (Republican Party) secures more electoral votes (Trump: 268 vs. Harris: 259), winning key states such as Texas, Florida, and Ohio. Although Kamala Harris (Democratic Party) is expected to perform well in traditional Democratic strongholds like California, New York, and Illinois, the overall projection favors the Republican candidate. The results also highlight competitive dynamics in swing states such as Pennsylvania, Georgia, and Arizona, where slight shifts could influence the final outcome. However, if we adopt the 2020 election results for the excluded states (CT and RI), the final tally would be Trump: 268 vs. Harris: 270 (259 + 11), resulting in a narrow victory for Harris (Democratic Party).
@@
4 Prediction for the 2024 US Election
4.1 Experiment Settings
We extend our multi-step reasoning pipeline to predict the outcome of the 2024 US Presidential Election, featuring Donald Trump and Kamala Harris as the primary candidates, which thus introduce differences from the 2020 results. As in the 2020 simulations, we incorporate only the candidates’ policy positions and backgrounds as time-sensitive inputs, excluding real-time data from social media or news sources to maintain consistency and control within the model’s input space. We follow the same simulation setup used in §3.3, utilizing identical hardware and LLM APIs. Since GPT-4o’s training corpus extends only until October 2023, this experiment evaluates the model’s ability to make predictions on future, unseen data. After the election, these predictions can be compared with the actual results, offering a unique opportunity to assess the LLM’s predictive capabilities.
Consistent with our prior methods, we simulate individual votes cast exclusively for the Democratic and Republican parties, omitting No-Preference votes from the analysis. The preference for Trump (Republican Party) in each state is calculated using Eq. (3). We applied the V3 pipeline to 49 states, excluding Rhode Island (RI) and Connecticut (CT) due to missing data. As in previous experiments, we employed random sampling with proportional extraction of synthetic personas for each state, ensuring statistical significance while optimizing computational efficiency. This state-level analysis tests the adaptability of our multi-step pipeline to unseen election data and provides a scalable framework for forecasting election outcomes with synthetic personas.
4.2 Prediction Results
The 2024 forecast by the LLM reveals several notable shifts compared to the 2020 prediction. A key change is in Wisconsin (WI), where Trump is now projected to win with 54.90% of the vote, marking a significant shift from 2020. Additionally, Trump shows improved performance in other swing and tipping-point states, including Pennsylvania (PA) with 47.85%, Michigan (MI) with 48.87%, and New Hampshire (NH) with 46.39%. However, despite these gains, he is still forecasted to narrowly lose these states to Harris.
For other critical states, the LLM predicts Trump will carry Arizona (AZ) with 51.09% of the vote, flipping it back to the Republicans, but lose North Carolina (NC) with 45.69%, consistent with the 2020 forecast despite the incorrect actual result in that election. In deep red states such as Texas (TX) with 56.36% and Alabama (AL) with 67.20%, Trump retains strong support. Similarly, Kamala Harris is expected to maintain dominance in traditional Democratic strongholds like California (CA) with 19.18%, New York (NY) with 22.41%, and Illinois (IL) with 30.36%.
The overall results indicate competitive dynamics in key swing states, where narrow margins may determine the final outcome. As shown in Figure 6, Trump is projected to win several battleground states like Georgia (GA) and Florida (FL) with 44.36% and 53.62%, respectively, strengthening his position. However, Harris still holds the edge in certain critical regions like Nevada (NV) with 34.77%.
Prediction Results. Based on the LLM’s predictions and the winner-takes-all electoral vote allocation, Trump is projected to win 268 electoral votes compared to Harris’s 259, securing a narrow victory. However, if we were to include the excluded states—Connecticut (CT) and Rhode Island (RI)—using their 2020 results, Harris’s total would rise to 270 electoral votes, resulting in a razor-thin margin in favor of the Democratic candidate.
Limitations. The 2024 prediction pipeline relies solely on synthetic persona demographics, the policy positions of both parties, and the candidates’ biographical information. While this controlled setup ensures consistency in inputs, it introduces several limitations. First, the model does not account for shifts in public opinion, media narratives, or unexpected events—such as economic shocks or political scandals—that could significantly impact voter behavior. Second, the simulation assumes static demographics between 2020 and 2024, meaning any population shifts, demographic changes, or state-level migration patterns (even if relatively small) are not factored into the prediction. As such, the objective of this paper is not to deliver definitive forecasts but rather to evaluate the LLM’s ability to generalize on unseen data and explore how well these generalizations align with real-world outcomes. This experimental approach provides a novel framework for assessing the model’s predictive capabilities and adaptability in dynamic political environments, offering valuable insights into the potential—and limitations—of LLMs for election forecasting.

### v2

# Towards More Accurate US Presidential Election via Multi-step Reasoning with Large Language Models

@@ Abstract
Can Large Language Models (LLMs) accurately predict election outcomes? While LLMs have demonstrated impressive performance in various domains, including healthcare, legal analysis, and creative tasks, their ability to forecast elections remains unknown. Election prediction poses unique challenges, such as limited voter-level data, rapidly changing political landscapes, and the need to model complex human behavior. To address these challenges, we introduce a multi-step reasoning framework designed for political analysis. Our approach is validated on real-world data from the American National Election Studies (ANES) 2016 and 2020, as well as synthetic personas generated by the leading machine learning framework, offering scalable datasets for voter behavior modeling. To capture temporal dynamics, we incorporate candidates’ policy positions and biographical details, ensuring that the model adapts to evolving political contexts. Drawing on Chain of Thought prompting, our multi-step reasoning pipeline systematically integrates demographic, ideological, and time-dependent factors, enhancing the model’s predictive power.
Towards More Accurate US Presidential Election via Multi-step Reasoning with Large Language Models
Chenxiao Yu1, Zhaotian Weng1, Yuangang Li1, Zheng Li2, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Carnegie Mellon University {cyu96374, wengzhao, yuangang, yzhao010}@usc.edu, winston@arimadata.com, xiyanghu@cmu.edu Corresponding authors
Important Notice Ongoing Work: This research is ongoing as of November 27, 2024. Research Integrity: This research is conducted independently w/o any funding. Content Warning: This paper may contain some offensive content generated by LLMs. Disclaimer: This study explores LLMs’ capacity in election forecasting. Predictions do not reflect the authors’ views and should not be interpreted as definitive forecasts.